# ZSMerge: Zero-Shot KV Cache Compression for Memory-Efficient Long-Context LLMs

### v2

@@ Abstract
The linear growth of key-value (KV) cache memory and quadratic computational complexity in attention mechanisms pose significant bottlenecks for large language models (LLMs) in long-context processing. While existing KV cache optimization methods address these challenges through token pruning or feature merging, they often incur irreversible information loss or require costly parameter retraining. To this end, we propose ZSMerge, a dynamic KV cache compression framework designed for efficient cache management, featuring three key operations: (1) fine-grained memory allocation guided by multi-dimensional token importance metrics at head-level granularity, (2) a residual merging mechanism that preserves critical context through compensated attention scoring, and (3) a zero-shot adaptation mechanism compatible with diverse LLM architectures without requiring retraining. ZSMerge significantly enhances memory efficiency and inference speed with negligible performance degradation across LLMs. When applied to LLaMA2-7B, it demonstrates a 20:1 compression ratio for key-value cache retention (reducing memory footprint to 5% of baseline) while sustaining comparable generation quality, coupled with triple throughput gains at extreme 54k-token contexts that eliminate out-of-memory failures. The code is available at https://github.com/SusCom-Lab/ZSMerge.
Machine Learning, ICML
@@ 1 Introduction
The advancement of the Transformer architecture has revolutionized sequence data processing, with Large Language Models (LLMs) being a prime illustration. LLMs have achieved remarkable progress, surpassing human capabilities in diverse applications. While multi-turn dialogue and long-context understanding are critical scenarios, handling extended sequences remains challenging due to the linear expansion of the Key-Value (KV) cache (which stores intermediate attention keys and values during generation to avoid redundant computations) in proportion to text length. For instance, deploying a LLaMA2-7B model will consume about 26GB of GPU memory; when the token length reaches 32K, the KV cache alone occupies 32G of memory, becoming the primary memory consumption. This, along with the quadratic computational complexity, significantly restricts LLM integration and performance, thus motivating recent research efforts toward optimizing KV cache utilization for enhanced inference efficiency.
To mitigate the large memory footprint of the KV cache, various optimization strategies have emerged. Some modify the core model architecture (Sarah et al., 2024; Ashkboos et al., 2024) or employ quantization techniques for lower precision representation (Dettmers et al., 2022; Xiao et al., 2023a). Other methods directly target the KV cache bottleneck during inference using context-aware techniques. Common among these is sparse approximation or token eviction, retaining only crucial tokens while discarding others (Zhang et al., 2024b, a); however, this can disrupt KV embeddings and cause information loss. To mitigate this, alternative context-aware techniques merge or fuse information from similar tokens (Dong et al., 2024), often using low-rank compression, though these may introduce architectural overhead or require model retraining.
@@
  • Fine-grained memory allocation. ZSMerge makes full use of the historical contribution of tokens, spatio-temporal characteristics, and its own data distribution characteristics to conduct fine-grained, head-level memory management. This adaptability allows it to maintain superior generation quality compared to the sparse strategies and achieve performance comparable to retraining-dependent approaches, even when operating with significantly reduced memory budgets. This ensures effectiveness across diverse long-context tasks.
  • Zero-shot and low-cost Integration. Designed for ease of use, ZSMerge does not introduce any model parameters and thus requires no retraining or model fine-tuning to adapt to different compression ratios or various task types. This also makes it compatible to various mainstream LLM architectures such as LLaMA, Falcon, and Mistral, thereby minimizing deployment overhead.
  • Efficient and throughput-boosting strategy. Characterized by linear operational complexity, ZSMerge’s compression mechanism is computationally lightweight. This efficiency brings significant throughput improvement, e.g., it helps achieve over triple the inference throughput compared to the original model at a token length of 54K. Furthermore, we have proved that the KV cache compressed by ZSMerge can effectively preserves the informational contribution of retained tokens, preventing signal degradation even as context length increases substantially.
@@ 2 Related Work
Figure 1: Comparative analysis of KV cache optimization strategies: (A) Sparse approximation directly addresses memory-compute growth by evicting low-contribution tokens, albeit inducing attention distribution drift. (B) Token merge mitigates sparse caching errors, often by learning residuals implicitly or explicitly, yet introduces overhead from auxiliary networks or fine-tuning. (C) ZSMerge (proposed) achieves zero-shot compression competitive with methods requiring learned components, doing so without added parameters while demonstrating superior generalization.
The escalating memory and computational demands of large language models (LLMs), particularly from KV cache growth in long-context scenarios, have spurred diverse optimization strategies. These are broadly categorized into context-agnostic and contextually adaptive approaches.
@@ 2.1 Context-Agnostic Optimization
Context-agnostic methods universally reduce resource consumption by altering the model’s structure or numerical representation, irrespective of the input sequence. Foundational techniques include Structural Compression via knowledge distillation (e.g., MiniLLM (Gu et al., 2024)) or matrix factorization (e.g., SliceGPT (Ashkboos et al., 2024)). Attention Mechanism Optimization, such as head pruning (Voita et al., 2019) and key-value sharing in Multi-Query (MQA (Shazeer, 2019)) and Grouped-Query Attention (GQA (Ainslie et al., 2023)), significantly cuts memory bandwidth. Simultaneously, Numerical Optimization tackles precision redundancy through techniques like post-training quantization (e.g., GPTQ (Frantar et al., 2022)), hardware-aligned quantization (e.g., INT-FlashAttention (Chen et al., 2024)), and exploiting low-rank properties (e.g., Linformer (Wang et al., 2020)). These strategies collectively alleviate memory bottlenecks via architectural simplification and precision reduction.
@@ 2.2 Contextually Adaptive Optimization
While context-agnostic methods offer baseline efficiency, recent research increasingly aims to directly target the KV cache—a critical bottleneck in long-sequence tasks driven by storage growth and quadratic computational complexity. The observed sparsity in attention values (Ji et al., 2021; Deng et al., 2024) motivates dynamic, context-aware strategies. Figure 1 illustrates the main categories of these adaptive approaches:
@@
(A) Sparse Approximation (Eviction): This class directly addresses memory-compute growth by evicting low-contribution tokens, based on the premise that not all history is equally vital. Work like (Xiao et al., 2023b) identified ”attention sinks,” leading to methods like StreamingLLM (Xiao et al., 2023b) which retains initial tokens alongside a sliding window. LM-infinite (Han et al., 2024) uses specialized masks to balance local and global context. H2O (Zhang et al., 2024b) employs attention score thresholds for dynamic eviction, while PyramidKV (Zhang et al., 2024a) uses layer-wise attention patterns for progressive caching. Dynamic routing (Roy et al., 2021) also fits here. However, eviction strategies inevitably induce attention distribution drift and error propagation by permanently discarding tokens (Dao et al., 2022; Xiao et al., 2023b).
@@
(B) Token Merging: To mitigate information loss from eviction, token merging fuses similar tokens instead of discarding them. Pioneered in vision (e.g., ToMe (Bolya et al., 2022)), it applies to text as well. Dynamic Memory Compression adaptively merges spatially proximate tokens, often requiring fine-tuning. LESS (Dong et al., 2024) uses auxiliary networks to compress tokens with minimal attention impact. While merging mitigates sparse caching errors by learning residuals (implicitly or explicitly), it often introduces overhead from auxiliary networks or fine-tuning, complicating deployment.
@@
(C) ZSMerge (Proposed): Addressing the limitations of both prior approaches, we propose ZSMerge. It achieves zero-shot compression competitive with sparse+residual methods without added parameters, while demonstrating superior generalization across task domains and compression ratios. ZSMerge offers a parameter-free, tuning-free approach, balancing efficiency and generation performance.
@@ 3.1 Preliminaries
$\mathbf{K}_{T}=[\mathbf{k}_{1},\mathbf{k}_{2},\ldots,\mathbf{k}_{T}]^{\top}\in\mathbb{R}^{T\times d},$ (1)
$\quad\mathbf{V}_{T}=[\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{T}]^{\top}\in\mathbb{R}^{T\times d},$
@@
$a_{t}^{(T)}$ $=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}{\sum_{s=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{s}/\sqrt{d})},$ (2)
$\quad\mathbf{o}^{(T)}$ $=\sum_{t=1}^{T}a_{t}^{(T)}\mathbf{v}_{t}$
@@ 3.2 Budget Allocation
where $B_{p}$, $B_{c}$, and $B_{r}$ govern context preservation, residual absorption, and proximity maintenance, respectively. The compressed cache matrices aggregate three complementary components:
@@
[…]
The proximity component $(\mathbf{K}_{p},\mathbf{V}_{p})$ preserves the latest $B_{p}$ tokens, capturing local attention patterns. The context component $(\mathbf{K}_{c},\mathbf{V}_{c})$ retains top-$B_{c}$ tokens ranked by contribution scores $\mathbf{s}^{(T)}\in\mathbb{R}^{T}$, which quantify contextual saliency. The residual component $(\mathbf{K}_{r},\mathbf{V}_{r})$ dynamically merges $B_{r}$ historically evicted tokens through learned transformations. This configuration subsumes eviction-based methods when $B_{r}=0$, while enabling continuous interpolation between sparse and dense representations through budget reallocation.
@@ 3.3 Contribution Evaluation
where the decay factor $\lambda\in[0,1]$ operates as a temporal discounting parameter analogous to reinforcement learning credit assignment, governing the exponential decay of historical attention contributions. We recommend setting it to a relatively large value to balance the contributions of long-term and short-term attention scores. Throughout this paper, $\lambda$ is configured as $0.98$ unless otherwise specified, as scores from 300 steps earlier become attenuated to approximately 0.002 through exponential decay ($0.98^{300}\approx 0.002$).
@@ 3.4 Residual Token Merging
    Update the selected slot through momentum-based fusion:
@@
Figure 2: KV cache operation process with $B_{r}=2$, $B_{c}=4$, $B_{p}=3$, and sequence length expansion $T=12\rightarrow 17$: residual merging (2 slots), context preservation (top-4) and proximity maintenance (latest 3).
@@
15:              Merge token $(\mathbf{k}_{\hat{c}},\mathbf{v}_{\hat{c}})$ into $\mathbf{K}_{r}$ via (6)-(7)
@@
3.5 Attention Output Stabilization
Algorithm 1 progressively merges low-saliency tokens through dynamic cache updates. To faithfully reconstruct the original attention distribution from compressed representations, we introduce a compensated attention scoring mechanism. The revised attention computation evolves from Eq. 2 to:
$\hat{a}_{t}^{(T)}$ $=\frac{\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d}+\alpha\log w_{t}\right)}{\sum_{i=1}^{T}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d}+\alpha\log w_{i}\right)},$ (8)
$\mathbf{\hat{o}}^{(T)}$ $=\sum_{t=1}^{T}\hat{a}_{t}^{(T)}\mathbf{v}_{t},$
where $w_{i}$ represents the fusion count of token $i$ (with $w_{i}=1$ for uncompressed tokens). The hyperparameter $\alpha\in(0,1]$ regulates residual compensation intensity, with empirical analysis suggesting an optimal range of $\alpha\in[0.6,0.8]$ for balanced performance. Notably, setting $\alpha=0$ reduces the merging algorithm to a degenerate case equivalent to a pure eviction policy, eliminating residual compensation entirely.
@@ 3.5 Attention Output Stabilization
Proof.
$\leq w_{r}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r}/\sqrt{d}\right)$
where the first inequality uses $\alpha\leq 1$, and the second applies Jensen’s inequality to the convex exponential function. For uncompressed tokens ($w_{i}=1$), we derive:
@@ 4 Experimental Evaluations
In this chapter, we aim to verify the effectiveness of our proposed method in terms of performance preservation and memory reduction. Firstly, we introduce the experimental setup, including the selected model architectures, the baseline methods for comparison, and parameters such as the datasets used. Secondly, we verify the performances of our method on memory control and throughput. Moreover, we compare the generation quality of different methods in the field of long texts and tasks. Finally, we discuss the trade-off between memory, time, and performance as well as the impact of relevant hyperparameters on the model.

### v3

@@ Abstract
The linear growth of key-value (KV) cache memory and quadratic computational complexity in attention mechanisms pose significant bottlenecks for large language models (LLMs) in long-context processing. While existing KV cache optimization methods address these challenges through token pruning or feature merging, they often incur irreversible information loss or require costly retraining. To this end, we propose ZSMerge, a dynamic KV cache compression framework for efficient cache management, featuring three key operations: (1) fine-grained memory allocation guided by multi-dimensional token importance metrics at head-level granularity, (2) a residual merging mechanism that preserves critical context through compensated attention scoring, and (3) a zero-shot adaptation mechanism compatible with diverse LLM architectures without retraining. ZSMerge significantly enhances memory efficiency and inference speed with negligible performance degradation. When applied to LLaMA2-7B, it achieves a 20:1 compression ratio for KV cache retention (reducing memory footprint to 5% of baseline) while sustaining comparable generation quality and tripling throughput at extreme 54k-token contexts, eliminating out-of-memory failures. These gains extend across diverse LLM families and long-context benchmarks, with ZSMerge consistently surpassing state-of-the-art cache optimization methods across tasks such as summarization, QA, reasoning, and code completion. The code is available at https://github.com/SusCom-Lab/ZSMerge.
@@ 1 Introduction
The advancement of the Transformer architecture has revolutionized sequence data processing, with Large Language Models (LLMs) being a prime illustration (vaswani2017attention, ). LLMs have achieved remarkable progress, surpassing human performance in diverse applications. While multi-turn dialogue and long-context understanding are critical scenarios (yi2024survey, ; gao2023self, ; li2023loogle, ; liu2024lost, ), handling extended sequences remains challenging due to the linear expansion of the Key-Value (KV) cache, which stores intermediate attention keys and values during generation to avoid redundant computations (kwon2023efficient, ). For instance, deploying a LLaMA2-7B (sarah2024llama, ) model will consume about 26GB of GPU memory; when the token length reaches 32K, the KV cache alone occupies 32GB of memory, becoming the primary memory consumption. This, along with the quadratic computational complexity, significantly restricts LLM integration and performance, thus motivating recent research efforts toward optimizing KV cache utilization for enhanced inference efficiency (zhang2024unifying, ; sun2024shadowkv, ; wang2024squeezeattention, ).
To mitigate the large memory footprint of the KV cache, various optimization strategies have emerged. Some modify the core model architecture (sarah2024llama, ; ashkboos2024slicegpt, ) or employ quantization techniques for lower precision representation (dettmers2022gpt3, ; xiao2023smoothquant, ). Other methods directly target the KV cache bottleneck during inference using context-aware techniques. Common among these is sparse approximation or token eviction, retaining only crucial tokens while discarding others (zhang2024h2o, ; zhang2024pyramidkv, ); however, this can disrupt KV embeddings and cause information loss. To mitigate this, alternative context-aware techniques merge or fuse information from similar tokens (dong_get_2024, ), often using low-rank compression, though these may introduce architectural overhead or require model retraining.
@@
  • Fine-grained memory allocation: ZSMerge makes full use of the historical contribution of tokens, spatio-temporal characteristics, and intrinsic data distribution characteristics to conduct fine-grained, head-level memory management. This adaptability allows it to maintain superior generation quality compared to the sparse strategies and achieve performance comparable to retraining-dependent approaches, even when operating with significantly reduced memory budgets. This ensures effectiveness across diverse long-context tasks.
  • Zero-shot and low-cost Integration: Designed for ease of use, ZSMerge does not introduce any model parameters and thus requires no retraining or model fine-tuning to adapt to different compression ratios or various task types. This also makes it compatible to various mainstream LLM architectures such as LLaMA, Falcon (almazrouei2023falcon, ), Mistral (jiang2023mistral, ), Qwen (yang2024qwen2, ), and Yi (aiYiOpenFoundation2025, ), thereby minimizing deployment overhead.
  • Efficient throughput-boosting strategy: Characterized by linear operational complexity, ZSMerge’s compression mechanism is computationally lightweight. This efficiency brings significant throughput improvement, e.g., it helps achieve over triple the inference throughput compared to the original model at a token length of 54K. Furthermore, we have proved that the KV cache compressed by ZSMerge can effectively preserves the informational contribution of retained tokens, preventing signal degradation even as context length increases substantially.
@@ 2 Related Work
KV cache growth in long-context LLMs creates severe memory and computational bottlenecks, motivating optimization strategies across two categories: context-agnostic and contextually adaptive approaches.
@@ 2.1 Context-Agnostic Optimization
Context-agnostic methods reduce resource consumption through structural or numerical modifications independent of input sequences. Key approaches include: (1) Structural compression via knowledge distillation (gu2024minillm, ) or matrix factorization (ashkboos2024slicegpt, ); (2) Attention optimization through head pruning (voita2019analyzing, ) and key-value sharing in Multi-Query (shazeer2019fast, ) and Grouped-Query Attention (ainslie2023gqa, ); (3) Numerical optimization via post-training quantization (frantar2022gptq, ), hardware-aligned quantization (chen2024int, ), and low-rank approximation (wang2020linformer, ).
@@ 2.2 Contextually Adaptive Optimization
Context-aware methods directly target KV cache bottlenecks by exploiting attention sparsity (ji2021distribution, ; deng2024attention, ). Figure 1 illustrates three main approaches:
Figure 1: Comparative analysis of contextually adaptive KV cache optimization strategies. (A) Sparse Approximation (Eviction): Low-contribution tokens are permanently discarded, reducing memory but risking error propagation. (B) Token Merging: Similar tokens are fused to mitigate information loss, but may require fine-tuning or auxiliary networks. (C) ZSMerge (Proposed): Achieves zero-shot "sparse+residual" compression without added parameters or tuning, balancing efficiency and performance.
@@
Sparse Approximation (Eviction): This class of methods directly tackles the growth of memory and computation by evicting low-contribution tokens, under the assumption that not all historical context is equally important. For example, StreamingLLM xiao2023efficient identifies "attention sinks" and retains both the initial tokens and a sliding context window. LM-infinite han2024lm uses specialized masks to balance local and global context. SnapKV li2024snapkv performs one-time KV cache pruning during the prefilling stage based on relatively consistent attention patterns, but lacks dynamic adjustment during decoding. H2O zhang2024h2o employs attention score thresholds for dynamic eviction, while PyramidKV zhang2024pyramidkv uses layer-wise attention patterns for progressive caching. Recent advances include FlexPrefill laiFlexPrefillContextAwareSparse2025 , which applies context-aware sparse attention specifically during prefilling. Dynamic routing roy2021efficient also fits here. However, eviction strategies inevitably induce attention distribution drift and error propagation by permanently discarding tokens dao2022flashattention ; xiao2023efficient .
@@
Token Merging: To mitigate information loss from eviction, token merging fuses similar tokens instead of discarding them. Pioneered in vision (e.g., ToMe bolya2022token ), it applies to text as well. Dynamic Memory Compression nawrot2024dynamic adaptively merges spatially proximate tokens, often requiring fine-tuning. LESS dong_get_2024 uses auxiliary networks to compress tokens with minimal attention impact. ThinK xuThinKThinnerKey2024 takes a different approach by applying query-driven pruning at the channel level, orthogonal to token-level compression. While merging mitigates sparse caching errors by learning residuals (implicitly or explicitly), it often introduces overhead from auxiliary networks or fine-tuning, complicating deployment.
Training-Dependent and System-Level Approaches: Recent work has explored complementary directions to token-level compression. Training-dependent methods include NSA (Native Sparse Attention) yuanNativeSparseAttention2025 , which introduces MLP-based compressor modules trained jointly with the attention mechanism, and MoBA (Mixture of Block Attention) luMoBAMixtureBlock2025 , which performs block-level pruning via gating with auxiliary routing networks. System-level approaches like OmniKV haoOmniKVDynamicContext2024 focus on dynamic context selection and offloading strategies for efficient long-context processing. InfiniPot kimInfiniPotInfiniteContext2024 adopts periodic block-wise compression with a fixed KV cache budget to reduce runtime overhead. While powerful, these methods require retraining, architectural modifications, or specialized system configurations.
@@
ZSMerge (Proposed): Addressing the limitations of prior approaches, we propose ZSMerge. It achieves zero-shot compression competitive using the "sparse+residual" method without added parameters, while demonstrating superior generalization across task domains and compression ratios. ZSMerge offers a parameter-free, tuning-free approach that is complementary to training-dependent and system-level methods, balancing efficiency and generation performance.
@@ 3.1 Preliminaries
$\mathbf{K}_{T}=[\mathbf{k}_{1},\mathbf{k}_{2},\ldots,\mathbf{k}_{T}]^{\top}\in\mathbb{R}^{T\times d},~\mathbf{V}_{T}=[\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{T}]^{\top}\in\mathbb{R}^{T\times d},$ (1)
@@
$a_{t}^{(T)}=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}{\sum_{i=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d})},~\mathbf{o}^{(T)}=\sum_{t=1}^{T}a_{t}^{(T)}\mathbf{v}_{t}.$ (2)
@@ 3.2 Budget Allocation
where $B_{p}$, $B_{c}$, and $B_{r}$ govern proximity maintenance, context preservation, and residual absorption, respectively. This tripartite division enables a nuanced approach to compression. Our empirical findings (see Appendix C.2) indicate that allocating the largest share of the budget to $B_{c}$ and $B_{p}$ is generally effective, as it prioritizes the retention of core information and local context. In contrast, $B_{r}$ is typically assigned a smaller budget to handle fine-grained residual adjustments. The compressed cache matrices therefore integrate three complementary components:
@@
with $\|$ denoting row-wise concatenation.
The proximity component $(\mathbf{K}_{p},\mathbf{V}_{p})$ preserves the latest $B_{p}$ tokens, capturing local attention patterns. The context component $(\mathbf{K}_{c},\mathbf{V}_{c})$ retains top-$B_{c}$ tokens ranked by contribution scores $\mathbf{s}^{(T)}\in\mathbb{R}^{T}$, which quantify contextual saliency. The residual component $(\mathbf{K}_{r},\mathbf{V}_{r})$ dynamically merges $B_{r}$ historically evicted tokens through attention-weighted aggregation transformations. The residual component $(\mathbf{K}_{r},\mathbf{V}_{r})$ maintains $B_{r}$ dedicated token slots, representing a compressed summary of previously evicted tokens. During generation, the residual cache is updated dynamically, progressively merging newly expelled tokens into the existing compressed representation. This configuration subsumes eviction-based methods when $B_{r}=0$.
@@ 3.3 Contribution Evaluation
where the decay factor $\lambda\in[0,1]$ acts as a temporal discounting parameter analogous to reinforcement learning credit assignment, controlling the exponential decay of historical attention contributions.
In practice, we find that the performance is not highly sensitive to the precise value of $\lambda$. We therefore fix $\lambda=0.98$ throughout this paper for simplicity, which yields a smooth balance between long-term and short-term contributions.
@@ 3.4 Residual Token Merging
    Update the selected slot via incremental mean aggregation:
@@
3.5 Attention Output Stabilization
@@
[…]
@@
Figure 2: KV cache operation process
Algorithm 1 progressively merges low-saliency tokens through dynamic cache updates. To more accurately reconstruct the attention distribution from compressed representations, we introduce a compensated attention scoring mechanism. The revised attention computation evolves from Eq. 2 to:
$\hat{a}_{t}^{(T)}=\frac{\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d}+\alpha\log w_{t}\right)}{\sum_{i=1}^{T}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d}+\alpha\log w_{i}\right)},~\mathbf{\hat{o}}^{(T)}=\sum_{t=1}^{T}\hat{a}_{t}^{(T)}\mathbf{v}_{t},$ (8)
where $w_{i}$ represents the fusion count of token $i$ (with $w_{i}=1$ for uncompressed tokens). The hyperparameter $\alpha\in[0,1]$ provides a soft transition, with $\alpha=0$ degenerating to an eviction-like policy. Its concrete influence is demonstrated in the appendix, and we fix $\alpha=1$ throughout our experiments.
@@
The proof is provided in Appendix A.
@@ 4 Experimental Evaluations
In this section, we begin by outlining the experimental setup, including model architectures, datasets, and baseline methods. We then assess efficiency improvements in terms of memory and speed. Next, we provide a detailed numerical error analysis to highlight the role of representation bias correction. Finally, we evaluate generation quality under diverse cache budgets, across multiple task types, and over different base model series, while comparing against a broad set of baselines.